# Conditional Logical Message Passing Transformer for Complex Query Answering

### v1

@@ 4.4. Answering Complex Queries with CLMPT
5. Experiments
5.1. Experimental Settings
5.1.1. Datasets and Queries
We evaluate our method on three commonly-used standard knowledge graphs: FB15k (Bordes et al., 2013), FB15k-237 (Toutanova and Chen, 2015), and NELL995 (Xiong et al., 2017a). For a fair comparison with previous works, we use the datasets of complex queries proposed by BetaE (Ren and Leskovec, 2020), which consist of five conjunctive query types ($1p/2p/3p/2i/3i$), five query types with atomic negations ($2in/3in/inp/pin/pni$), and four zero-shot query types ($pi/ip/2u/up$), as illustrated in Figure 3. The datasets provided by BetaE introduce complex queries with hard answers that cannot be obtained by traversing the KG directly. That is, the evaluation focuses on discovering such hard answers to complex queries. We follow previous works (Ren and Leskovec, 2020; Wang et al., 2023b; Zhang et al., 2021) and train our model with five conjunctive query types and five query types with atomic negations. When evaluating the model, we consider all query types including the four zero-shot query types. The statistics for each dataset are presented in Table 1. More details about the datasets can be found in Appendix A.
@@ 5.1.1. Datasets and Queries
5.1.2. Evaluation Protocol
5.1.3. Baselines
We consider the state-of-the-art neural complex query answering models for EFO-1 queries in recent years as our baselines to compare: BetaE (Ren and Leskovec, 2020), ConE (Zhang et al., 2021), Q2P (Bai et al., 2022), MLP (Amayuelas et al., 2022), GammaE (Yang et al., 2022), CylE (Nguyen et al., 2023), CQD-CO (Arakelyan et al., 2020), and LMPNN (Wang et al., 2023b), where CQD-CO and LMPNN use pre-trained neural link predictor, while other models do not. In addition, some other neural CQA models cannot handle logical negation operators, such as BIQE (Kotnis et al., 2021), PREM (Choudhary et al., 2021a), kgTransformer (Liu et al., 2022), etc., which use the EPFO queries generated by Q2B (Ren* et al., 2020) for training and evaluation. We compare these works on the Q2B datasets in Appendix C. We also compare our work with symbolic integration CQA models in Appendix E.
5.1.4. Model Details
Following LMPNN (Wang et al., 2023b) and CQD-CO (Arakelyan et al., 2020), we choose ComplEx-N3 (Lacroix et al., 2018; Trouillon et al., 2016) as the neural link predictor in our work and use the ComplEx-N3 checkpoints released by CQD-CO to conduct experiments. The rank of ComplEx-N3 is 1,000, and the epoch for the checkpoints is 100. For all datasets, we use a two-layer transformer encoder with eight self-attention heads, where the dimension of the hidden layer of the feed-forward network (FFN) in the transformer is 8,192. For the training objective, the negative sample size $K$ is 128, and $T$ is chosen as 0.05 for FB15k-237 and FB15k and 0.1 for NELL995. For more details about the implementation and experiments, please refer to Appendix B.
@@ 5.1.4. Model Details
5.2. Main Results
@@ 5.3.1. Conditional Logical Meassage Passing
5.3.2. Hyperparamaters and Pooling Approaches
5.3.3. Frozen or Trainable Neural Link Predictors
@@
Although CLMPT, which uses the trainable neural link predictor, has more training parameters, it still significantly outperforms LMPNN on EPFO queries across all datasets when the neural link predictor is frozen. This reflects the effectiveness of CLMPT. For more experimental analyses of the model parameters, please refer to Appendix D.
@@ 6. Conclusion
In this paper, we propose CLMPT, a special message passing neural network, to answer complex queries over KGs. Based on one-hop inference by pre-trained neural link predictor on atomic formulas, CLMPT performs logical message passing conditionally on the node type. In the ablation study, we verify that this conditional message passing mechanism can effectively reduce the computational costs and even improve the performance of the model to some extent. Furthermore, CLMPT uses the transformer to aggregate messages and update the corresponding node embedding. Through the self-attention mechanism, CLMPT can explicitly model logical dependencies between messages received by a node and between messages and that node by assigning adaptive weights to the messages and the node. The experimental results show that CLMPT achieves a strong performance through this transformer-based node embedding update scheme. Future work may integrate symbolic information into CLMPT to improve the performance in answering negative queries.
Appendix A More Details about The Datasets
The statistics of the three knowledge graphs used in our experiment are shown in Table 6. As we described in section 5.1.1, we conduct experiments on the complex query datasets proposed in the BetaE paper (Ren and Leskovec, 2020). The statistics for the complex queries contained in these datasets are shown in Table 1. The statistics for the answers for each query type are shown in Table 7.
As mentioned in Section 5.1.3, there are some other neural CQA models (Liu et al., 2022; Choudhary et al., 2021a; Kotnis et al., 2021) that are trained and evaluated on the EPFO query datasets proposed in the Q2B paper (Ren* et al., 2020). To compare these models, we also conduct experiments (see Appendix C) on the FB15k-237 and NELL995 configured by the Q2B setting. We denote these two datasets as FB237-Q2B and NELL-Q2B. The statistics for the EPFO queries and corresponding answers for Q2B datasets are shown in Table 8 and Table 9.
One difference between the BetaE datasets and the Q2B datasets is the limit on the number of answers to the queries, since the Q2B datasets have queries with more than 5000 answers, almost $1/3$ of some datasets. Furthermore, the Q2B datasets do not contain queries with negation. These differences mean that BetaE (Ren and Leskovec, 2020) datasets are more challenging, so most CQA models in recent years chose to evaluate their performance on the BetaE datasets. To obtain convincing comparison results, we mainly conduct experiments on the BetaE datasets to compare existing competitive neural CQA models.
Table 8. The number of training, validation, and test queries generated for different query structures of the Q2B datasets (Ren* et al., 2020).
Queries Training Validation Test
Dataset 1p others 1p Others 1p Others
FB237-Q2B 149,689 149,689 20,101 5,000 22,812 5,000
NELL-Q2B 107,982 107,982 16,927 4,000 17,034 4,000
Table 9. The average number of answers divided by query structure for Q2B datasets (Ren* et al., 2020).
Dataset 1p 2p 3p 2i 3i ip pi 2u up
FB237-Q2B 13.3 131.4 215.3 69.0 48.9 593.8 257.7 35.6 127.7
NELL-Q2B 8.5 56.6 65.3 30.3 15.9 310.0 144.9 14.4 62.5
Appendix B More experimental details
@@
Our code is implemented using PyTorch. We use NVIDIA Tesla A10 GPU (24GB) and NVIDIA GeForce RTX 3090 GPU (24GB) to conduct all of our experiments. We use the AdamW (Loshchilov and Hutter, 2019) optimizer, whose weight decay is 1e-4, to tune the parameters. The learning rate is 5e-5 for NELL995 and FB15k-237 and 1e-4 for FB15k. The batch size is 512 for NELL995 and FB15k-237 and 1,024 for FB15k. The training epoch for all datasets is 100. Following CQD-CO (Arakelyan et al., 2020) and LMPNN (Wang et al., 2023b), we choose ComplEx-N3 (Lacroix et al., 2018; Trouillon et al., 2016) as the neural link predictor for fair comparison. In this case, we consider the logical message encoding function $\rho$ corresponding to Equations 1013. For these equations, the $\lambda$ is a hyperparameter that needs to be determined. In LMPNN application, for simplicity, LMPNN just lets $3\lambda\left\|\cdot\right\|=1$ and then all denominators in these closed-form expressions are 1, namely:
@@
Table 10. The Hits@3 results of other neural CQA baselines on answering EPFO queries generated by Q2B (Ren* et al., 2020). $\mathbf{Avg_{p}}$ represents the average results of these EPFO query types. The boldface indicates the best results.
Dataset Model 1p 2p 3p 2i 3i pi ip 2u up $\mathbf{Avg_{p}}$
@@
Appendix C Comparison with More neural CQA models on Q2B Datasets
To further evaluate the performance of CLMPT, we also consider comparing neural CQA models that cannot handle logical negation and are trained and evaluated on Q2B datasets, including GQE (Hamilton et al., 2018), Q2B (Ren* et al., 2020), BIQE (Kotnis et al., 2021), PERM (Choudhary et al., 2021a), kgTransformer (Liu et al., 2022), and SILR (Wang et al., 2023c). Among these models, BIQE, kgTransformer, and SILR all use the transformer (Vaswani et al., 2017) to encode complex queries directly. BIQE and SILR serialize the query graph and treat the CQA task as the sequence learning task to solve. kgTransformer uses the pre-training and fine-tuning strategies based on Heterogeneous Graph Transformer (HGT) (Hu et al., 2020) to answer complex queries. In essence, kgTransformer is also a transformer-based graph neural network. We use the results reported in these papers (Liu et al., 2022; Wang et al., 2023c) for comparison. Since they report the Hits@3 results, for a fair comparison, we also use Hits@3 as the evaluation metric, which calculates the proportion of correct answer entities ranked among the top 3.
@@ Appendix C Comparison with More neural CQA models on Q2B Datasets
Appendix D Analyses on Model Parameters
Appendix E Comparison with Symbolic Integration Models
@@
According to (Ren et al., 2023; Wang et al., 2023b), for larger knowledge graphs, neural CQA models and symbolic integration models have different scalabilities. As we discussed in Section 2, symbolic integration models require more computing resources and are more likely to suffer from scalability issues. For example, GNN-QE employs NBFNet (Zhu et al., 2021) to compute message passing on the whole KG, resulting in complexity that is linear to $(|\mathcal{E}|+|\mathcal{V}|)d$, where $|\mathcal{E}|$ is the number of edges in KG, $|\mathcal{V}|$ is the number of nodes in KG, and $d$ is the embedding dimension. For CLMPT, which only performs message passing on the query graph, the complexity is just $O(d)$. Besides, GNN-QE estimates the probability of whether each entity is the answer at each intermediate step, making the size of its fuzzy sets scale linearly with $|\mathcal{V}|$. As a result, GNN-QE requires much more computational costs, requiring 128GB GPU memory to run a batch size of 32. For CLMPT-small(F), which has only 32M trainable parameters, it only requires less than 12GB GPU memory to run a batch size of 1024. Even in this case, CLMPT-small(F) can still achieve performance competitive with GNN-QE on NELL995.
We suspect that integrating symbolic information into CLMPT can improve the performance, especially on negative queries. However, exploring how to integrate symbolic information into CLMPT is beyond the scope of this paper. This extension is left for future work.

### v2

@@ 4.4. Answering Complex Queries with CLMPT
5. Experiments
5.1. Experimental Settings
5.1.1. Datasets and Queries
We evaluate our method on three commonly-used standard knowledge graphs: FB15k (Bordes et al., 2013), FB15k-237 (Toutanova and Chen, 2015), and NELL995 (Xiong et al., 2017a). For a fair comparison with previous works, we use the datasets of complex queries proposed by BetaE (Ren and Leskovec, 2020), which consist of five conjunctive query types ($1p/2p/3p/2i/3i$), five query types with atomic negations ($2in/3in/inp/pin/pni$), and four zero-shot query types ($pi/ip/2u/up$), as illustrated in Figure 3. The datasets provided by BetaE introduce complex queries with hard answers that cannot be obtained by traversing the KG directly. That is, the evaluation focuses on discovering such hard answers to complex queries. We follow previous works (Ren and Leskovec, 2020; Wang et al., 2023b; Zhang et al., 2021) and train our model with five conjunctive query types and five query types with atomic negations. When evaluating the model, we consider all query types including the four zero-shot query types. The statistics for each dataset are presented in Table 1.
@@ 5.1.1. Datasets and Queries
5.1.2. Evaluation Protocol
5.1.3. Baselines
We consider the state-of-the-art neural complex query answering models for EFO-1 queries in recent years as our baselines to compare: BetaE (Ren and Leskovec, 2020), ConE (Zhang et al., 2021), Q2P (Bai et al., 2022), MLP (Amayuelas et al., 2022), GammaE (Yang et al., 2022), CylE (Nguyen et al., 2023), CQD-CO (Arakelyan et al., 2020), and LMPNN (Wang et al., 2023b), where CQD-CO and LMPNN use pre-trained neural link predictor, while other models do not. In addition, some other neural CQA models cannot handle logical negation operators, such as BIQE (Kotnis et al., 2021), PREM (Choudhary et al., 2021a), kgTransformer (Liu et al., 2022), etc., which use the EPFO queries generated by Q2B (Ren* et al., 2020) for training and evaluation. We compare these works on the Q2B datasets in Appendix B. We also compare our work with symbolic integration CQA models in Appendix C.
5.1.4. Model Details
Following LMPNN (Wang et al., 2023b) and CQD-CO (Arakelyan et al., 2020), we choose ComplEx-N3 (Lacroix et al., 2018; Trouillon et al., 2016) as the neural link predictor in our work and use the ComplEx-N3 checkpoints released by CQD-CO to conduct experiments. The rank of ComplEx-N3 is 1,000, and the epoch for the checkpoints is 100. For all datasets, we use a two-layer transformer encoder with eight self-attention heads, where the dimension of the hidden layer of the feed-forward network (FFN) in the transformer is 8,192. For the training objective, the negative sample size $K$ is 128, and $T$ is chosen as 0.05 for FB15k-237 and FB15k and 0.1 for NELL995. For more details about the implementation and experiments, please refer to Appendix A.
@@ 5.3.1. Conditional Logical Meassage Passing
5.3.2. Hyperparamaters and Pooling Approaches
5.3.3. Frozen or Trainable Neural Link Predictors
@@
Although CLMPT, which uses the trainable neural link predictor, has more training parameters, it still significantly outperforms LMPNN on EPFO queries across all datasets when the neural link predictor is frozen. This reflects the effectiveness of CLMPT.
5.4. Analyses on Model Parameters
@@ 6. Conclusion
In this paper, we propose CLMPT, a special message passing neural network, to answer complex queries over KGs. Based on one-hop inference by pre-trained neural link predictor on atomic formulas, CLMPT performs logical message passing conditionally on the node type. In the ablation study, we verify that this conditional message passing mechanism can effectively reduce the computational costs and even improve the performance of the model to some extent. Furthermore, CLMPT uses the transformer to aggregate messages and update the corresponding node embedding. Through the self-attention mechanism, CLMPT can explicitly model logical dependencies between messages received by a node and between messages and that node by assigning adaptive weights to the messages and the node. The experimental results show that CLMPT achieves a strong performance through this transformer-based node embedding update scheme. Future work may integrate symbolic information into CLMPT to improve the performance in answering negative queries. For limitations, please refer to Appendix D.
Acknowledgements.
We thank the anonymous reviewers for their helpful feedbacks. The work described in this paper was partially funded by the National Natural Science Foundation of China (Grant Nos. 62272173, 62273109), the Natural Science Foundation of Guangdong Province (Grant Nos. 2024A1515010089, 2022A1515010179), the Science and Technology Planning Project of Guangdong Province (Grant No. 2023A0505050106), and the National Key R&D Program of China (Grant No. 2023YFA1011601).
Appendix A More Experimental Details
@@
Our code is implemented using PyTorch. We use NVIDIA Tesla A10 GPU (24GB) and NVIDIA GeForce RTX 3090 GPU (24GB) to conduct all of our experiments. We use the AdamW (Loshchilov and Hutter, 2019) optimizer, whose weight decay is 1e-4, to tune the parameters. The learning rate is 5e-5 for NELL995 and FB15k-237 and 1e-4 for FB15k. The batch size is 512 for NELL995 and FB15k-237 and 1,024 for FB15k. The training epoch for all datasets is 100. Following CQD-CO (Arakelyan et al., 2020) and LMPNN (Wang et al., 2023b), we choose ComplEx-N3 (Lacroix et al., 2018; Trouillon et al., 2016) as the neural link predictor for fair comparison. In this case, we consider the logical message encoding function $\rho$ corresponding to Equations 1013. For these equations, $\lambda$ is a hyperparameter that needs to be determined. In LMPNN application, for simplicity, LMPNN just lets $3\lambda\left\|\cdot\right\|=1$ and then all denominators in these closed-form expressions are 1, namely:
@@
Table 7. The Hits@3 results of other neural CQA baselines on answering EPFO queries generated by Q2B (Ren* et al., 2020). Avg represents the average results of all query types. The boldface indicates the best results.
Dataset Model 1p 2p 3p 2i 3i pi ip 2u up Avg
@@
Appendix B Comparison with More Neural CQA Models on Q2B Datasets
To further evaluate the performance of CLMPT, we also consider comparing neural CQA models that cannot handle logical negation and are trained and evaluated on Q2B datasets (Ren* et al., 2020), including GQE (Hamilton et al., 2018), Q2B (Ren* et al., 2020), BIQE (Kotnis et al., 2021), PERM (Choudhary et al., 2021a), kgTransformer (Liu et al., 2022), and SILR (Wang et al., 2023c). Among these models, BIQE, kgTransformer, and SILR all use the transformer (Vaswani et al., 2017) to encode complex queries directly. BIQE and SILR serialize the query graph and treat the CQA task as the sequence learning task to solve. kgTransformer uses the pre-training and fine-tuning strategies based on Heterogeneous Graph Transformer (HGT) (Hu et al., 2020) to answer complex queries. In essence, kgTransformer is also a transformer-based graph neural network. We use the results reported in these papers (Liu et al., 2022; Wang et al., 2023c) for comparison. Since they report the Hits@3 results, for a fair comparison, we also use Hits@3 as the evaluation metric, which calculates the proportion of correct answer entities ranked among the top 3.
Appendix C Comparison with Symbolic Integration Models
@@
According to (Ren et al., 2023; Wang et al., 2023b), for larger knowledge graphs, neural CQA models and symbolic integration models have different scalabilities. As we discussed in Section 2, symbolic integration models require more computing resources and are more likely to suffer from scalability issues. For example, GNN-QE employs NBFNet (Zhu et al., 2021) to compute message passing on the whole KG, resulting in complexity that is linear to $(|\mathcal{E}|+|\mathcal{V}|)d$, where $|\mathcal{E}|$ is the number of edges in KG, $|\mathcal{V}|$ is the number of nodes in KG, and $d$ is the embedding dimension. For CLMPT, which only performs message passing on the query graph, the complexity is just $O(d)$. Besides, GNN-QE estimates the probability of whether each entity is the answer at each intermediate step, making the size of its fuzzy sets scale linearly with $|\mathcal{V}|$. As a result, GNN-QE requires much more computational costs, requiring 128GB GPU memory to run a batch size of 32. For CLMPT-small(F), which has only 32M trainable parameters, it only requires less than 12GB GPU memory to run a batch size of 1024. Even in this case, CLMPT-small(F) can still achieve performance competitive with GNN-QE on NELL995. We suspect that integrating symbolic information into CLMPT can improve the performance, especially on negative queries. However, exploring how to integrate symbolic information into CLMPT is beyond the scope of this paper. This extension is left for future work.
Appendix D Limitations
Due to the use of the transformer architecture in the node embedding update scheme in CLMPT, more computational costs are required compared to some previous works based on multi-layer perceptron. In addition, CLMPT only achieves sub-optimal performance on negative queries. We suspect that integrating symbolic information into CLMPT may improve the performance on negative queries.